Tracy is thinking of a 3-digit number less that 700. If she switches the ones and the hundreds digit, the number increases by 396. If she instead switches the tens ones the ones digits, the number decreased by 18. What is the original number?
Answer by tmwilliamlin168(27)

x*100+y*10+z<700
z*100+y*10+x=x*100+y*10+z+396
x*100+z*10+y=x*100+y*10+z-18
the answer is 397 due to my calculations in my head
substitute it if u don't believe me